1. 4a+5[1-3(2a-1)]=2(2a+5)

Answer by venugopalramana(3286)

4a+5[1-3(2a-1)]=2(2a+5)
4a+5-5*3(2a-1)=2*2a+2*5
4a+5-15*2a+15=4a+10
4a-30a-4a=10-5-15=-10
-30a=-10
a=-10/-30=1/3